# Reactive Diffusion Policy: Slow-Fast Visual-Tactile Policy Learning for Contact-Rich Manipulation

### v1

@@ VII Conclusion
In this paper, we presented TactAR, a low-cost teleoperation system providing real-time tactile / force feedback through AR, and Reactive Diffusion Policy (RDP), a novel slow-fast imitation learning algorithm for contact-rich manipulation. TactAR demonstrates that high-quality tactile /force feedback can be achieved through AR visualization without expensive specialized hardware. RDP successfully addresses the trade-off between sequence modeling and closed-loop control through its hierarchical design - using a slow network for complex trajectory planning and a fast network for reactive tactile feedback control. Through extensive experiments on three challenging contact-rich tasks, we demonstrated that RDP significantly outperforms state-of-the-art visual IL baselines in terms of both task completion and reactivity to tactile feedback. The cross-sensor experiments further validated that our approach generalizes well across different tactile / force sensors. We believe this work takes an important step toward making visual-tactile imitation learning more practical and accessible.

### v2

@@ VII Conclusion
In this paper, we present TactAR, a low-cost teleoperation system that provides real-time tactile / force feedback through AR, and Reactive Diffusion Policy (RDP), a novel slow-fast imitation learning algorithm for contact-rich manipulation. TactAR demonstrates that high-quality tactile / force feedback can be achieved through AR visualization without expensive specialized hardware. RDP successfully addresses the trade-off between sequence modeling and closed-loop control through its hierarchical design - using a slow network for complex trajectory planning and a fast network for reactive tactile feedback control. Through extensive experiments on three challenging contact-rich tasks, we demonstrated that RDP significantly outperforms state-of-the-art visual IL baselines in terms of both task completion and reactivity to tactile feedback. The cross-sensor experiments further validated that our approach generalizes well across different tactile / force sensors. We believe that this work takes an important step toward making visual-tactile imitation learning more practical and accessible.
Acknowledgments
We thank Yongkai Fan for providing inspiration for the hardware design in our experiments. We thank Yi Wang, Junjie Fang, Fangyuan Zhou, Yuan Fang, Yanbing Zhou, Zimo Wen, Yutong Li, Wei Jiang, Yijin Chen and Yongkai Fan for participating in the user study. We also thank Hongjie Fang and Zihao He for providing help and feedback on manipulation with force sensing.